OpenStudy (fixxer):

Help me find a transfor for the followin inequality to be a rectangle which can be solved by using one single double integral

OpenStudy (fixxer):

\[0\le x+y \le \pi \]\[3\le 2x.y \le 7\]

OpenStudy (fixxer):

2x-y sorry for that typo

OpenStudy (irishboy123):

do the Jacobian :) i think, as you have typo's, that you mean a co-ordinate shift like this: \(u = x + y, \qquad v = 2x -y \) with \(u \in [0, \pi], v \in [3,7]\) and you have the transform \(dx~ dy \to |J| ~ du ~ dv\)

OpenStudy (fixxer):

Hi, Thank you very much for the transform.

OpenStudy (irishboy123):

don't thank me mate, thank whoever invented partial differentiation :)

OpenStudy (fixxer):

lol :p they aint alive anymore I guess, anyway since we are at it is the jacobian the determinant of this matrix??\[\left[\begin{matrix}\frac{ 1 }{ 3 } & \frac{ 1 }{ 3 } \\ \frac{ 2 }{ 3 } & -\frac{ 1 }{ 3 }\end{matrix}\right]\] Which will give me -1/3 and trapped area of -4?

OpenStudy (irishboy123):

i'm getting -1/3 too for \(\det \left[\begin{matrix}1 & 1 \\ 2 & -1\end{matrix}\right]\)

OpenStudy (fixxer):

Great! means im doing at least a small part right lols